OpenStudy (anonymous):

The reciprocal of sqrt(2) is half of what number?

OpenStudy (anonymous):

The reciprocal of \(\sqrt{2}\) is \(\frac{1}{\sqrt{2}}=\frac{\sqrt{2}}{2}\). So the reciprocal is half of \(\sqrt{2}\).